Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-04-15

Night report 2017-04-15

Observers: Claudia Gutiérrez, Regis Cartier
TAT on duty: Joe Anderson, Kate Maguire
Support team: Francesco Taddia, Jesper Sollerman, Christoffer Fremling

Daytime calibrations: PESSTO_EFOSC_bias, PESSTO_EFOSC_calib 

Weather conditions
Completely cloudy at the beginning of the night (The dome was closed until 00:00 UT)
At 00:00 UT: Sky partly cloudy, seeing of 1.2", humidity ~ 20%, wind ~ 6m/s. The observations started at 00:15 UT.  
At 04:00 UT: Sky completely covered (Dome closed)
No more observations during the night due to the clouds.

Seeing: 1.2 arcsec (00:21 UT), 1 arcsec (03:18 UT)

UT        Target            RA                     DEC             Time spent(min)   slit/grism or filter         Description
00:17    AT2017cuh    05:34:52.24    -20:02:48.10   27 + 8                    gr13                            Classification: SN Ia at ~7 days after max, z=0.053. 
00:55    SN2017cyy    09:36:36.30     -63:56:54.6    37 + 3                    gr11                            Follow-up. SN Ia few days before max.
01:36    SN2017cyy    09:36:36.30    -63:56:54.6     40                          gr16                            Follow-up. SN Ia few days before max.      
02:20    AT2017cfo     00:38:12.91    +28:07:02.0    40 + 7                    gr13                            Follow-up. SN II, Halpha with a weak absorption component.     
03:09    AT2017awz    11:07:35.48    +22:51:04.5    35                          gr13                            Follow-up. SN Ia with decent signal.   
03:45    LTT3864        10:32:13 .90    -35:37:42.4    15                          g11+g16                     Standard telluric --Thanks to the TO!!! :) --



Comments